hey im kushal and i need to find all the combinations of 6 numbers that are obtained from 1 to 40 only.
please reply me as some as possible before 14 nov 2009- 17:00

Answer by Theo(2973)

numbers are 1 to 40

you draw 6 numbers.

each can be 1-40.

you are allowed duplicates.

possible combinations would be:

40^6 = 4096000000

to see how this works, use smaller numbers.

suppose you could pick 2 number from 1 to 2 each.

total number of combinations would be 2^3 = 8

2^3 means 2 raised to the power of 3 means 2*2*2

* means to multiply (2*2*2 = 8)

they would be:

111
112
121
122
211
212
221
222

suppose you could pick 3 numbers from 1 to 3 each.

total combinations would be 3^3 = 27

they would be:

111
112
113
121
122
123
131
132
133
211
212
213
221
222
223
231
232
233
311
312
313
321
322
323
331
332
333

same concept applies with 6 number from 1 to 40 each.

answer is 40^6